# Singularities of symmetric powers and irrationality of motivic zeta functions

### v1

@@ Abstract.
Let $K_{0}(\mathcal{V}_{K})$ be the Grothendieck ring of varieties over a field $K$ of characteristic zero, and let $\mathbb{L}=[\mathbb{A}^{1}_{K}]$ denote the Lefschetz class. We prove that if a $K$-variety has $\mathbb{L}$-rational singularities, then all its symmetric powers also have $\mathbb{L}$-rational singularities. We then use this result to show that, for a smooth complex projective variety $X$ of dimension greater than one, the rationality of its Kapranov motivic zeta function $Z(X,t)$ (viewed as a formal power series over $K_{0}(\mathcal{V}_{K})$) implies that the Kodaira dimension of $X$ is negative and that $X$ does not admit global nonzero differential forms of even degree. This extends the irrationality part of the Larsen–Lunts rationality criterion from the surface case to arbitrary dimension. We also discuss some applications of these results.
@@ 1. Introduction
This result tells us that for a smooth connected projective variety $Y$, the image of its class $[Y]$ in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$ records the stable birational equivalence class of $Y$. However, if $Y$ is not smooth, then the geometric interpretation of its class in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$ is less clear.
For a quasi-projective variety $X$ over $K$, we denote by $\operatorname{Sym}^{d}(X)$ the $d$-fold symmetric power of $X$. By definition, it is the quotient $X^{d}/S_{d}$ by the permutation action of the symmetric group $S_{d}$ on $X^{d}$. We adopt the convention that $\operatorname{Sym}^{0}(X)=\operatorname{Spec}{K}$.
The ring $K_{0}(\mathcal{V}_{K})$ has a natural $\lambda$-structure (in the sense of Definition 2.8), defined by $\lambda^{n}([X])=[\operatorname{Sym}^{n}(X)]$ [^2] As pointed out in [LL2, Section 8], this $\lambda$-structure is not special.. The Kapranov motivic zeta function of $X$ is defined as the formal power series
@@
M. Kapranov conjectured that it is natural to expect that $Z(X,t)$ is a rational function (as a formal power series over $K_{0}(\mathcal{V}_{K})$) for an arbitrary variety $X$  [Ka, Remark 1.3.5]. He showed that this is the case if $X$ is a curve with a rational point, and D. Litt later generalized this result to arbitrary geometrically connected curves (see  [Litt, Theorem 10]). However, it turned out that rationality fails in dimension two. It was proved by M. Larsen and V. Lunts  [LL2, Theorem 1.1] that for a smooth complex projective surface $X$ the zeta function $Z(X,t)$ is rational if and only if $X$ has negative Kodaira dimension. In this paper we prove the following generalization of the irrationality part of this result to higher dimensions:
@@
Let $X$ be a smooth complex projective variety of dimension greater than one. If $Z(X,t)$ is rational as a formal power series with coefficients in $K_{0}(\mathcal{V}_{\mathbb{C}})$, then the Kodaira dimension of $X$ is negative and $H^{0}(X,\Omega^{2i}_{X})=0$ for all $i>0$.
@@
    in $K_{0}(\mathcal{V}_{K})/({\mathbb{L}})$?
@@
We use the notion of $\mathbb{L}$-rational singularities (Def. 2.1), introduced in  [NS, Definition 4.2.4]. As proved in [NS, Lemma 4.2.5], if $Y$ has $\mathbb{L}$-rational singularities, then $\widetilde{Y}]=[Y]$ in $K_{0}(\mathcal{V}_{K})/({\mathbb{L}})$ for any resolution of singularities $\widetilde{Y}$ of $Y$. We prove that if a quasi-projective variety $X$ has $\mathbb{L}$-rational singularities, then so do all of its symmetric powers (Theorem 2.12). This result, which may be of independent interest, in particular implies an affirmative answer to 1.3(a). We first establish a certain lemma about the Grothendieck ring of varieties (Lemma 2.10), which allows us to use an inductive argument on $n$ to prove that $\operatorname{Sym}^{n}(\mathbb{A}^{k})$ has $\mathbb{L}$-rational singularities for all $n,k\geqslant 0$ (Proposition 2.15). We then show that the general case of Theorem 2.12 can be deduced from this.
We then prove Theorem 1.2 in Section 3 (see Theorem 3.3 for a more precise formulation). Our proof largely relies on the ideas of [LL2], though we generalize their strategy in several ways. As mentioned above, our main generalization is that we work with an arbitrary resolution $\widetilde{\operatorname{Sym}}^{m}(X)$. This is made possible mainly by Theorem 2.12 and the result of D.Arapura and S. Archava, which states that singularities of ${\operatorname{Sym}}^{m}(X)$ are canonical  [AA, Proposition 1]. More precisely, Theorem 2.12 implies that $\mu_{k}(\operatorname{Sym}^{m}(X))=\mu_{k}(\widetilde{\operatorname{Sym}}^{m}(X))$ for all $k\geqslant 1$, while  [AA, Proposition 1] helps us study $\mu_{k}(\widetilde{\operatorname{Sym}}^{m}(X))$. We then show that, under the assumptions of Theorem 3.3, one can find $k$ such that $\mu_{k}(Z(X,t))$ is not rational.
Finally, in Section 4, we discuss a couple of applications of the results from Sections 2 and 3. We compute the image of $Z(B,t)$ in $(K_{0}(\mathcal{V}_{K})/(\mathbb{L}))[[t]]$ for a Severi-Brauer variety $B$ (Proposition 4.1) and generalize a result of D. Litt ([Litt2, Theorem 19]) on stable birationality of symmetric powers (Proposition 4.3). We also note that the positive answer to Question 1.3(a) implies that the motivic measure $\mu_{1}:K_{0}(\mathcal{V}_{K})\to\mathbb{Z}[M]$ is a homomorphism of $\lambda$-rings (Proposition 4.4) which was already observed by Larsen and Lunts.
Let us also mention that, for the purposes of motivic integration, it is more natural to consider the localized ring $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$ instead of $K_{0}(\mathcal{V}_{K})$. Thus, one might ask when $Z(X,t)$ is rational as a formal power series over $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$. We do not investigate this question in this paper, although we expect that the rationality of $Z(X,t)$ over $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$ should also impose strong restrictions on the geometry of $X$.
@@ 2. Singularities of symmetric powers
$\boxed{2.\Rightarrow 3.}$ Clear.
@@
A $\lambda$-ring is a ring with a $\lambda$-structure. A $\lambda$-homomorphism is a homomorphism of rings commuting with all $\lambda$-operations. It is clear that for every $\lambda$-ring $A$ the map
@@
is an additive group homomorphism [^3] For every ring $A$, one can also equip $1+tA[[t]]$ with a canonical $\lambda$-structure. If $A$ is in addition a $\lambda$-ring, then one says that $A$ is a special $\lambda$-ring if $\lambda_{t}$ is a $\lambda$-homomorphism. We will not need this notion in the rest of the article..
@@
It is clear that there are unique maps
@@
Theorem 2.12.
Remark 2.13.
It follows from [Ko2, Theorem 3.36] that over a field of characteristic zero there exists a functorial resolution of singularities, i. e. a functor $F:X\mapsto(\pi_{X}:\widetilde{X}\to X)$ from the category of varieties over $K$ and smooth morphisms to the category of smooth varieties and smooth morphisms such that $\pi_{X}:\widetilde{X}\to X$ is a resolution of singularities, and for any smooth $f:X\to Y$ the canonical map $\widetilde{X}\to X\times_{Y}\widetilde{Y}$ is an isomorphism. From now on in this section we assume that we have fixed a functorial resolution of singularities.
Notation 2.14.
@@
We will use the notation $\lambda=(1^{l_{1}},\ldots,n^{l_{n}})$ for a partition of $n$ which corresponds to a decomposition $n=d_{1}+\cdots+d_{r}$ such that the first $l_{1}$ of the $d_{i}$ are equal to $1$, the next $l_{2}$ of the $d_{i}$ are equal to $2$ and so on. Let $S_{\lambda}=S_{\lambda}(\mathbb{A}^{k})$ be as in Notation 2.14, and let $H_{\lambda}=\{g\in S_{n}\mid gS_{\lambda}=S_{\lambda}\}$ be the normalizer of $S_{\lambda}$. We have $H_{\lambda}=H_{1}\rtimes H_{2}$, where $H_{1}=\prod_{i=1}^{r}S_{d_{i}}$ and $H_{2}=\prod_{j}S_{l_{j}}$, where the second product is over all $j\in\{1,\ldots,n\}$ such that $l_{j}\neq 0$. Note that since $H_{1}$ acts trivially on $S_{\lambda}$, we have $\mathbb{A}^{k}_{\lambda}\cong S_{\lambda}/H_{\lambda}\cong(S_{\lambda}/H_{1})/H_{2}\cong S_{\lambda}/H_{2}$.
@@
Let $N=\{(x_{1},\ldots x_{n})\in(\mathbb{A}^{k})^{n}\mid x_{1}+\cdots+x_{n}=0\}\subset\mathbb{A}^{kn}$ be a linear subspace complementary to the small diagonal $S_{(n^{1})}\subset\mathbb{A}^{kn}$. It is preserved by the action of $S_{n}$ and we have an $S_{n}$-equivariant isomorphism $N\times S_{(n^{1})}\overset{\sim}{\longrightarrow}(\mathbb{A}^{k})^{n}$ given by $(x,y)\mapsto x+y$. Since the action of $S_{n}$ on $S_{(n^{1})}$ is trivial, it induces an isomorphism $N/S_{n}\times\Delta^{n}\overset{\sim}{\longrightarrow}\operatorname{Sym}^{n}(\mathbb{A}^{k})$. Since $\Delta^{n}_{\mathbb{A}^{k}}\cong\mathbb{A}^{k}$ is smooth, we have a Cartesian square
@@
For every point $x\in S_{\lambda}$ we have $\operatorname{Stab}_{S_{n}}(x)\subset H_{\lambda}$, so the quotient morphism $\mu:\mathbb{A}^{kn}/H_{\lambda}\to\operatorname{Sym}^{n}(\mathbb{A}^{k})$ is étale at $y$ for every point $y\in S_{\lambda}/H_{\lambda}$ by [SGA1, Exposé V. Proposition 2.2]. Therefore, there exist an open subset $W^{\prime}\subset\mathbb{A}^{kn}/H_{\lambda}$ containing $S_{\lambda}/H_{\lambda}$, such that $\left.\mu\right|_{W^{\prime}}$ is étale. Let $W$ be the preimage of $W^{\prime}$ under the projection $\mathbb{A}^{kn}\to\mathbb{A}^{kn}/H_{\lambda}$, so that $W^{\prime}=W/H_{\lambda}$. Consider a commutative diagram
$\prod_{i=1}^{n-1}\widetilde{\operatorname{Sym}}^{l_{i}}(\widetilde{\operatorname{Sym}}^{i}(\mathbb{A}^{k}))$$\widetilde{{W}/{H_{\lambda}}}$$\widetilde{\operatorname{Sym}}^{n}(\mathbb{A}^{k})$$\prod_{i=1}^{n-1}\operatorname{Sym}^{l_{i}}(\operatorname{Sym}^{i}(\mathbb{A}^{k}))$${W}/{H_{\lambda}}$$\operatorname{Sym}^{n}(\mathbb{A}^{k})$${S_{\lambda}}/{H_{\lambda}}$${S_{\lambda}}/{H_{\lambda}}$$\mathbb{A}^{k}_{\lambda}$$\scriptstyle{\beta\circ\alpha}$$\scriptstyle{\widetilde{\mu}}$$\scriptstyle{\pi_{W/H_{\lambda}}}$$\scriptstyle{\rho_{n}}$$\scriptstyle{\mu}\scriptstyle{\sim}\scriptstyle{\sim}$
where $\pi_{W/H_{\lambda}}$ is a functorial resolution. Since $\mu$ is étale and an isomorphism on $S_{\lambda}/H_{\lambda}$, it follows that all squares of the diagram are Cartesian and $\rho_{n}^{-1}(\mathbb{A}^{k}_{\lambda})\cong\pi^{-1}_{W/H_{\lambda}}(S_{\lambda}/H_{\lambda})$.

### v2

@@ Abstract.
Let $K_{0}(\mathcal{V}_{K})$ be the Grothendieck ring of varieties over a field $K$ of characteristic zero, and let $\mathbb{L}=[\mathbb{A}^{1}_{K}]$ denote the Lefschetz class. We prove that if a $K$-variety has $\mathbb{L}$-rational singularities, then all its symmetric powers also have $\mathbb{L}$-rational singularities. We then use this result to show that, for a smooth complex projective variety $X$ of dimension greater than one, the rationality of its Kapranov motivic zeta function $Z(X,t)$ (viewed as a formal power series over $K_{0}(\mathcal{V}_{\mathbb{C}})$) implies that the Kodaira dimension of $X$ is negative and that $X$ does not admit global nonzero differential forms of even degree. This extends the irrationality part of the Larsen–Lunts rationality criterion from the surface case to arbitrary dimension. We also discuss some applications of these results.
@@ 1. Introduction
Theorem 1.1 tells us that for a smooth connected projective variety $Y$, the image of its class $[Y]$ in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$ records the stable birational equivalence class of $Y$. However, if $Y$ is not smooth, then the geometric interpretation of its class in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$ is less clear.
For a quasi-projective variety $X$ over $K$, we denote by $\operatorname{Sym}^{n}(X)$ the $n$-fold symmetric power of $X$. By definition, it is the quotient of $X^{n}$ by the permutation action of the symmetric group $S_{n}$. We adopt the convention that $\operatorname{Sym}^{0}(X)=\operatorname{Spec}{K}$.
The ring $K_{0}(\mathcal{V}_{K})$ has a natural $\lambda$-structure (in the sense of Def. 2.8), defined by $\lambda^{n}([X])=[\operatorname{Sym}^{n}(X)]$ [^2] As pointed out in [LL2, Section 8], this $\lambda$-structure is not special.. The Kapranov motivic zeta function of $X$ is defined as the formal power series
@@
M. Kapranov conjectured that it is natural to expect that $Z(X,t)$ is a rational function (as a formal power series over $K_{0}(\mathcal{V}_{K})$) for an arbitrary variety $X$  [Ka, Remark 1.3.5]. He showed that this is the case if $X$ is a curve with a rational point, and D. Litt later generalized this result to arbitrary geometrically connected curves  [Litt2, Theorem 10]. However, it turned out that rationality fails in dimension two. It was proved by M. Larsen and V. Lunts  [LL2, Theorem 1.1] that for a smooth complex projective surface $X$, the zeta function $Z(X,t)$ is rational if and only if $X$ has negative Kodaira dimension. Moreover, they proved irrationality in a stronger sense by showing that if $\kappa(X)\geqslant 0$, then $Z(X,t)$ is not pointwise rational (see Def. 3.1). In this paper we prove the following generalization of this result to higher dimensions:
@@
Let $X$ be a smooth connected complex projective variety of dimension $d>1$. Assume that at least one of the following conditions hold:
  • (a)
  • (b)
    There exists $i>0$ such that $H^{0}(X,\Omega_{X}^{2i})\neq 0$.
Then $Z(X,t)$ is not pointwise rational.
@@
    in $K_{0}(\mathcal{V}_{K})/({\mathbb{L}})$? In other words, is the class $\widetilde{\operatorname{Sym}}^{m}(X)]-[\operatorname{Sym}^{m}(X)]$ always divisible by $\mathbb{A}^{1}]$?
@@
We use the notion of $\mathbb{L}$-rational singularities (Def. 2.1), introduced in  [NS, Definition 4.2.4]. As proved in [NS, Lemma 4.2.5], if $Y$ has $\mathbb{L}$-rational singularities, then $\widetilde{Y}]=[Y]$ in $K_{0}(\mathcal{V}_{K})/({\mathbb{L}})$ for any resolution of singularities $\widetilde{Y}$ of $Y$. We prove the following
Theorem 1.4.
This result, which may be of independent interest, in particular implies an affirmative answer to 1.3(a). We first establish a certain lemma about the Grothendieck ring of varieties (Lemma 2.10), which allows us to use an inductive argument on $n$ to prove that $\operatorname{Sym}^{n}(\mathbb{A}^{k})$ has $\mathbb{L}$-rational singularities for all $n,k\geqslant 0$ (Proposition 2.13). We then show that the general case of Theorem 1.4 can be deduced from this.
After that, we prove Theorem 1.2 in Section 3. Our proof largely relies on the ideas of [LL2], though we generalize their strategy in several ways. As mentioned above, our main generalization is that we work with an arbitrary resolution $\widetilde{\operatorname{Sym}}^{m}(X)$. This is made possible mainly by Theorem 1.4 and the result of D.Arapura and S. Archava, which states that singularities of ${\operatorname{Sym}}^{m}(X)$ are canonical  [AA, Proposition 1]. More precisely, Theorem 1.4 implies that $\mu_{k}(\operatorname{Sym}^{m}(X))=\mu_{k}(\widetilde{\operatorname{Sym}}^{m}(X))$ for all $k\geqslant 1$, while  [AA, Proposition 1] helps us study $\mu_{k}(\widetilde{\operatorname{Sym}}^{m}(X))$. We then show that, under the assumptions of Theorem 1.2, one can find $k$ such that $\mu_{k}(Z(X,t))$ is not rational.
Finally, in Section 4, we discuss a couple of applications of the results from Sections 2 and 3. We show that Theorem 1.4 together with the result of J. Kollár allows us to easily compute the image of $Z(B,t)$ in $(K_{0}(\mathcal{V}_{K})/(\mathbb{L}))[[t]]$ for a Severi-Brauer variety $B$ (Proposition 4.1). We then explain that, if $X$ satisfies the assumptions of Theorem 1.2, one can use the measures $\mu_{k}$ to detect which symmetric powers of $X$ are not stably birational to each other (Proposition 4.3), which generalizes [Litt, Theorem 19]. We also note that a positive answer to Question 1.3(a) implies that the motivic measure $\mu_{1}:K_{0}(\mathcal{V}_{K})\to\mathbb{Z}[M]$ is a homomorphism of $\lambda$-rings (Proposition 4.4), which was already observed by Larsen and Lunts.
Let us also mention that, for the purposes of motivic integration, it is more natural to consider the localized ring $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$ instead of $K_{0}(\mathcal{V}_{K})$. Thus, one might ask when $Z(X,t)$ is rational as a formal power series over $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$. We do not investigate this interesting question in this paper, although we expect that the rationality of $Z(X,t)$ over $K_{0}(\mathcal{V}_{K})[1/\mathbb{L}]$ should also impose strong restrictions on the geometry of $X$.
@@ 2. Singularities of symmetric powers
$\boxed{2.\Rightarrow 3.}$ Clear (by functoriality of the Grothendieck ring).
@@
A $\lambda$-ring is a ring with a $\lambda$-structure. A $\lambda$-homomorphism is a homomorphism of rings commuting with all $\lambda$-operations. It is clear that $(\lambda^{i}:A\to A)_{i\geqslant 0}$ define a $\lambda$-structure on $A$ if and only if the map
@@
is a homomorphism of groups.
For every ring $A$, one can equip $1+tA[[t]]$ with a canonical $\lambda$-structure. If $A$ is in addition a $\lambda$-ring, then one says that $A$ is a special $\lambda$-ring if $\lambda_{t}$ is a $\lambda$-homomorphism.
@@
It is straightforward to check that there are unique maps
@@
Before proving Theorem 1.4, let us fix some notation. Recall that it follows from [Ko2, Theorem 3.36] that over a field of characteristic zero there exists a functorial resolution of singularities, i. e. a functor $F:X\mapsto(\pi_{X}:\widetilde{X}\to X)$ from the category of varieties over $K$ and smooth morphisms to the category of smooth varieties and smooth morphisms such that $\pi_{X}:\widetilde{X}\to X$ is a resolution of singularities, and for any smooth $f:X\to Y$ the canonical map $\widetilde{X}\to X\times_{Y}\widetilde{Y}$ is an isomorphism. From now on in this section we assume that we have fixed a functorial resolution of singularities.
Notation 2.12.
@@
We will use the notation $\lambda=(1^{l_{1}},\ldots,n^{l_{n}})$ for a partition of $n$ which corresponds to a decomposition $n=d_{1}+\cdots+d_{r}$ such that the first $l_{1}$ of the $d_{i}$ are equal to $1$, the next $l_{2}$ of the $d_{i}$ are equal to $2$ and so on. Let $S_{\lambda}=S_{\lambda}(\mathbb{A}^{k})$ be as in Notation 2.12, and let $H_{\lambda}=\{g\in S_{n}\mid gS_{\lambda}=S_{\lambda}\}$ be the normalizer of $S_{\lambda}$. We have $H_{\lambda}=H_{1}\rtimes H_{2}$, where $H_{1}=\prod_{i=1}^{r}S_{d_{i}}$ and $H_{2}=\prod_{j}S_{l_{j}}$, where the second product is over all $j\in\{1,\ldots,n\}$ such that $l_{j}\neq 0$. Note that since $H_{1}$ acts trivially on $S_{\lambda}$, we have $\mathbb{A}^{k}_{\lambda}\cong S_{\lambda}/H_{\lambda}\cong(S_{\lambda}/H_{1})/H_{2}\cong S_{\lambda}/H_{2}$.
@@
Let $N=\{(x_{1},\ldots x_{n})\in(\mathbb{A}^{k})^{n}\mid x_{1}+\cdots+x_{n}=0\}\subset\mathbb{A}^{kn}$ be a linear subspace complementary to the small diagonal $S_{(n^{1})}\subset\mathbb{A}^{kn}$. It is preserved by the action of $S_{n}$ and we have an $S_{n}$-equivariant isomorphism $N\times S_{(n^{1})}\overset{\sim}{\longrightarrow}(\mathbb{A}^{k})^{n}$ given by $(x,y)\mapsto x+y$. Since the action of $S_{n}$ on $S_{(n^{1})}$ is trivial, it induces an isomorphism $N/S_{n}\times\Delta^{n}_{\mathbb{A}^{k}}\overset{\sim}{\longrightarrow}\operatorname{Sym}^{n}(\mathbb{A}^{k})$. Since $\Delta^{n}_{\mathbb{A}^{k}}\cong\mathbb{A}^{k}$ is smooth, we have a Cartesian square
@@
For every point $x\in S_{\lambda}$ we have $\operatorname{Stab}_{S_{n}}(x)\subset H_{\lambda}$, so the quotient morphism $\mu:\mathbb{A}^{kn}/H_{\lambda}\to\operatorname{Sym}^{n}(\mathbb{A}^{k})$ is étale at $y$ for every point $y\in S_{\lambda}/H_{\lambda}$ by [SGA1, Exposé V. Proposition 2.2]. Therefore, there exist an open subset $W^{\prime}\subset\mathbb{A}^{kn}/H_{\lambda}$ containing $S_{\lambda}/H_{\lambda}$, such that $\left.\mu\right|_{W^{\prime}}$ is étale. Moreover, by shrinking $W^{\prime}$ if needed we can assume that $W^{\prime}\cap\mu^{-1}(\mathbb{A}^{k}_{\lambda})=S_{\lambda}/H_{\lambda}$.
Let $W$ be the preimage of $W^{\prime}$ under the projection $\mathbb{A}^{kn}\to\mathbb{A}^{kn}/H_{\lambda}$, so that $W^{\prime}=W/H_{\lambda}$. Consider a commutative diagram
$\prod_{i=1}^{n-1}\widetilde{\operatorname{Sym}}^{l_{i}}(\widetilde{\operatorname{Sym}}^{i}(\mathbb{A}^{k}))$$\widetilde{{W}/{H_{\lambda}}}$$\widetilde{\operatorname{Sym}}^{n}(\mathbb{A}^{k})$$\prod_{i=1}^{n-1}\operatorname{Sym}^{l_{i}}(\operatorname{Sym}^{i}(\mathbb{A}^{k}))$${W}/{H_{\lambda}}$$\operatorname{Sym}^{n}(\mathbb{A}^{k})$${S_{\lambda}}/{H_{\lambda}}$${S_{\lambda}}/{H_{\lambda}}$$\mathbb{A}^{k}_{\lambda}$$\scriptstyle{\beta\circ\alpha}$$\scriptstyle{\widetilde{\mu}}$$\scriptstyle{\pi_{W/H_{\lambda}}}$$\scriptstyle{\rho_{n}}$$\scriptstyle{\left.\mu\right|_{W^{\prime}}}\scriptstyle{\sim}\scriptstyle{\sim}$
where $\pi_{W/H_{\lambda}}$ is a functorial resolution. Since $\left.\mu\right|_{W^{\prime}}$ is étale and $(\left.\mu\right|_{W^{\prime}})^{-1}(\mathbb{A}^{k}_{\lambda})=S_{\lambda}/H_{\lambda}$, it follows that all squares of the diagram are Cartesian and $\rho_{n}^{-1}(\mathbb{A}^{k}_{\lambda})\cong\pi^{-1}_{W/H_{\lambda}}(S_{\lambda}/H_{\lambda})$.